A manufacturing manager collected the following information for the Production Department:

ObservationMachine hoursTotal Operating Costs
January8,000$60,000
February9,20066,000
March7,60061,000
April8,80064,000
May9,00065,000

Required: 

a. Determine the estimating cost function with machine-hours as the cost driver using the high-low method. 

b. If June's estimated machine-hours total 8,400, what is the estimated cost of the Production Department? 
Answer #1
a Total cost incurred
Machine hours Costs incurred
High level of activity                                    9,200 $66,000
Low level of activity                                    7,600 $60,000
Change                                    1,600 $6,000
Variable maintainance cost per machine hours =$6,000 / 1,600 =$3.75 per machine hours
Fixed cost element =$66,000 - (9,200*$3.75)
Fixed cost element =$66,000 - $34,500 =$31,500
b Estimated Total cost at 8,400 hours =$31,500 + (8,400 hours*$3.75)
Estimated Total cost at 8,400 hours =$31,500 + 31,500 =$63,000
Answer #3

(a)

Slope coefficient= (Cost at highest activity – cost at lowest activity)/(Highest activity - Lowest activity)

($66000 - $61000)/(9200 - 7600)

= $3.125 per machine-hour

Constant = $66000 - ($3.125 × 9200) = $37250

Machine-hour estimating equation = $37250 + $3.125X

(b) Estimated machine hours = 8400

Estimated cost = $37250 + $3.125X

        = $37250 + ($3.125 * 8400) = $63500
Answer #2


Solution: a) MHS Total operating cost 66000 Highest level of activity Lowest level of activity Difference 9200 7600 1600 6100